# Event-Based Simulation of Stochastic Memristive Devices for Neuromorphic Computing

### v2

@@ III Metal Oxide Memristor Modelling
In this section, we look at the application of the modelling framework to the modelling of resistive drift phenomenon under zero bias in titanium dioxide memristors. As such, we give an illustrative example of the proposed event-based modelling framework, detailing a readout equation, switching probabilities, and the volatility state variable for this setting.
III-A Physical Model
@@
Modelling of metal oxide Metal-Insulator-Metal (MIM), and more specifically - Titanium Dioxide MIM - has been previously modelled as a single conductive pathway of varying resistance [9], and other similar devices, such as tantalum oxide memristors have been modelled as exhibiting the dominance of a single filament or conductive region [15].
@@
Based on initial investigation of a drift dataset (see Section III-E below), we model the conductance of the device as linearly proportional to state variable $n$. We hypothesise that the mechanism underlying the switching of the device modulates the width of a conductive region within the active substrate, bridging a connection between the electrodes. Phenomena that could account for this behaviour include a single conductive filament that grows or reduces in width, or multiple conductive filaments that each contribute linearly to the conductance of the device. We observe that it is necessary to use a value of $n_{\text{extra}}$ and $m_{\text{extra}}$ much greater than the value of $N$. This suggests a thresholding effect, where the rate of switching is relatively constant in the switching region, under a constant voltage bias. This corresponds somewhat to the ideal behaviour for a memristor, which is why such titanium dioxide devices were presented as candidate memristive devices originally [9]. We assume a large $n_{\text{extra}}=10\times N$ such that the switching rates remain relatively constant across the considered range of resistance values.
@@ III-B Readout Equation
$R(n)=\frac{1}{g_{\text{diff}}\max{(n,n_{\text{thresh}})}+\frac{1}{R_{\text{high}}}}$ (4)
We can see from the equation that the behaviour of the model is thresholded, with the resistance remaining at the value $R_{\text{high}}$ until the value of $n$ surpasses the parameter $0\leq n_{\text{thresh}}\leq N$, upon which the conductance increases linearly in steps of $g_{\text{diff}}$. This is motivated by the conductive filament model described above.
@@
$n(R)=n_{\text{thresh}}+\frac{1}{g_{\text{step}}}\left(\frac{1}{R}-\frac{1}{g_{\text{parallel}}}\right)$ (5)
@@
III-C Switching Probability
To determine the rates of the Poisson processes, we make use of a Boltzmann occupation function that is dependent on the voltage bias and the temperature, as well as static parameters, rather than the Fermi-Dirac occupation function (logistic function) used to determine the switching probabilities in the GMSM. This is due to it being a good model of probabilistic behaviour in microscopic thermodynamic systems, modulated by energy barriers.
@@ III-C Switching Probability
with the parameters defined in Table I:
$E_{a}=qV_{a}$ Activation energy (the energy barrier height). This increases or decreases both the up and down switching rates.
$E_{\text{off}}=qV_{\text{off}}$ Energy offset that leads to a shift of the equilibrium point from equal numbers of states being in the high or low resistive state. This adjusts the balance of switching rates under zero bias. (Added, rather than subtracted, for $\lambda^{\prime}_{down}$).
@@
$T(t)$ Time-dependent temperature variable in degrees Kelvin. This may be influenced by the history of applied voltages and the device state history, which directly impact the rate of switching, as well as the equilibrium point when $E_{\text{off}}$ is not 0.
$\rho(t)$ Volatility state variable. Has a similar impact to the temperature in affecting the global switching rate (in fact, the temperature can be considered a volatility state variable as well).
@@
In this model, application of a voltage bias, $V(t)$, lowers or raises the activation energy barrier in one direction, simultaneously raising or lowering it in the other direction.
III-D Volatility
@@
In modelling the effects of Joule heating, we can assume that a proportion of the current flowing through the device (proportional to the square of the current flowing through the device) is converted to heat energy and raises the internal temperature. We can also assume that Newton’s law of cooling applies simultaneously, resulting in power dissipation and restoration of the internal temperature to the equilibrium (bath) temperature.
To introduce Joule heating as a form of volatility, one can define the rate of change of temperature in terms of a thermal time constant that is a function of a thermal resistance ($R_{th}$) and thermal capacitance ($C_{th}$) [16], dictating how easily heat energy in the form of temperature captures and is dissipated from the device, wit the thermal time constant denoted $\tau_{th}=C_{th}R_{th}$. The differential equation for the internal temperature of the device is given as:
@@
where $\tau_{\text{volatile}}$ is the volatility time constant, dictating how quickly the equilibrium value is attained (and therefore how slowly a non-zero volatility decays in the absence of an input signal); and $c_{\text{volatile}}$ is the volatility factor, which - when multiplied by the applied bias - yields the equilibrium volatility (dictating the volatility state variable’s leverage on the switching rates).
III-E Dataset
@@
III-E1 Quantisation
Figure 5 shows a histogram of the unquantised resistance values in the dataset, showing the range of interest to be the interval of approximately $[10^{3},10^{5}]$, where changes to the value of $n$ significantly affect the output resistance. As such, we choose the values of $N$, $n_{\text{thresh}}$, $g_{\text{step}}$, and $g_{\text{parallel}}$ accordingly. We set $g_{\text{step}}=10^{-7}$, $g_{\text{parallel}}=10^{-10}$, and $N=2n_{\text{thresh}}=20,000$.
@@ III-E1 Quantisation
$\mathcal{R}_{\text{val}}=$$\{R(n_{\text{thresh}}),R(n_{\text{thresh}}+1),$
$\cdots,R(N-1),R(N)\}$
where $\begin{cases}|\mathcal{R}_{\text{bound}}|&=N-n_{\text{thresh}}+2\\ |\mathcal{R}_{\text{val}}|&=N-n_{\text{thresh}}+1\end{cases}$
III-E2 Pairs
@@
III-F Model Fitting
III-F1 Calculation of the Equilibrium Point (Metastable Point)
@@
$n_{\text{eq}}$ $=\frac{N}{e^{\frac{qV_{\text{off}}}{k_{B}T}}+1}$ (11)
@@
$V_{\text{off}}$ $=\ln{\left(\frac{N-n_{\text{eq}}}{n_{\text{eq}}}\right)}\cdot\frac{k_{B}T}{q}\equiv\ln{\left(\frac{m_{\text{eq}}}{n_{\text{eq}}}\right)}\cdot V_{T}$ (12)
@@
III-F2 Linear Conductance Approximation
@@
Figure 6: The mean change in resistance for the dataset of pairs with $\Delta T=10000$. We show a moving average of the data with different window sizes (5 and 50), for a clearer view of the non-linear trend.
@@
We use a zero-bias model consisting of an approximately constant change in the conductance over the given time interval, that is independent of the initial resistance. This assumption is based on our own model, which fits a trend in the observed data to a reasonable standard. We use an iterative trust region fitting algorithm [19] to minimise the squared error associated with fitting the function to the predicted resistance change. The model is given as follows, with parameter $a$ the change in the state variable - fit to the data:
$\Delta R(R_{\text{init}})=\frac{1}{g_{\text{parallel}}-ag_{\text{diff}}}-R_{\text{init}}$ (13)
For the chosen timestep of $10,000$ seconds, we find a value of $a=3.62$ yields the best fit, for the given parameters $g_{\text{diff}}$, $g_{\text{parallel}}$, $N$, and $n_{\text{thresh}}$. This associated line of best fit for this linear conductance change model is shown in Figure 6. The moving averages are plotted to enable a clearer view of the general trend. Although their shape is clear, they do include bias associated with the (symmetric) averaging window size, with a larger window size tending to overestimate the mean change, as can be seen from the larger magnitude of the moving average plot for a window size of 50, compared to the model fit to the data. The moving average window of size 5 yields a line whose mean closely matches the model.
III-F3 Parameter Fitting and Constant Rate Approximation
@@
We would like to obtain parameters that allow the mean behaviour of the model to correspond to the mean predictions of the linear conductance model. Due to the difficulty in obtaining an analytic expression for the delay-conditional distribution, we use a constant rate approximation to obtain approximate mean values for this distribution over time. We assume that the up and down switching rates of the model remain relatively constant. In order to fit the model, we calculate the rate at a value of $n=n_{\text{thresh}}$ and make the assumption that the rate remains equal to this rate over the region of interest. Thus, this approximation is only valid if the quantities $\frac{n_{\text{thresh}}-a}{n_{\text{thresh}}}$ and $\frac{N-n_{\text{thresh}}+a}{N-n_{\text{thresh}}}$, approximating the percentage change in the up and down switching rates remain relatively close to 1.
@@
We choose the parameters $N$ and $n_{\text{thresh}}$ accordingly. As such, we require $n_{\text{thresh}}>>a$ and $N-n_{\text{thresh}}>>a$. We must also ensure that $N+1-n_{\text{thresh}}$, defining the number of unique resistance states according to Equation 4, is appropriately selected to give a sufficiently fine quantisation over the switching region for which we have data.
@@
For the given value of the parameter $a$, we can determine an appropriate value of $V_{a}$. We first select a value of $V_{\text{off}}$ to ensure an equilibrium point (see Section III-F1) that is far above the active switching region (in terms of resistance), to ensure that the assumption that the rate is constant over the considered range holds, since the equilibrium point at a particular bias sets the zero-crossing point for the rate equation, hence our constant rate assumption will break down as the value of $n$ approaches $n_{\text{eq}}$. In our case, this corresponds to setting an equilibrium point with a value of $n<<n_{\text{thresh}}$. We can see in Figure 7 how the choice of $V_{\text{off}}$ impacts $n_{\text{eq}}$; and hence, the constant rate assumption, and the fit to the linear conductance model. Assuming a constant rate and a given value of $V_{\text{off}}$, we can derive a value of $V_{a}$ according to the following equation:
$V_{a}=-\frac{k_{B}T}{q}\ln{\left(2\Delta Tn_{\text{thresh}}\sinh{\left(V_{\text{off}}\frac{q}{2k_{B}T}\right)}\right)}$ (14)
@@
In Figure 8, we see averaged zero-bias simulations from the event-based model, based on the fit parameters, showing the mean change in the resistance fitting to the linear conductance model, as well as the associated variance of the evolutions (which increases with the magnitude of the resistance).
Figure 8: The mean change in the resistance following fitting to the linear conductance model, with errorbars indicating the standard deviation of series, for simulations with length $\Delta T=10000$.
III-G Simulation
In addition to demonstrating the fit of the zero-bias mean values to those determined according to the linear conductance model, we demonstrate the capabilities of the model to simulate memristive behaviour through a series of simulated experiments for a titanium dioxide thin film memristive device. Due to the event-driven nature of the state transition model we conduct our simulations in this section and in Section IV in Python, rather than a circuit simulation language such as SPICE.
III-G1 Experimental Parameters
The parameters used for the experiments, chosen for $\Delta T=10000$, are summarised in Table II. The choices for these parameters are given in Sections III-E1 and III-F3.
@@ III-G1 Experimental Parameters
III-G2 Switching Experiments
We use the model, with the derived parameters, to generate example device data for simulated input signals. Our first experiments assume there is no volatile behaviour. In the second set of experiments, we assume Joule heating - modulated by the current flowing through the device - and associated restoration of the equilibrium temperature according to Newton’s law of cooling, as described in Section III-D.
@@ III-G2 Switching Experiments
The input signal parameters are given in Table III. We simulate series from starting resistances ranging between $10k\Omega$ and $100k\Omega$. We choose the parameters for the input signals in order to highlight model behaviour. In all the experiments, we see that the model is biased towards movement to a higher resistance state, with negative pulses having a lesser impact on reducing the resistive state than positive pulses have on increasing the state. During the inactive period of each pulse, we see the device’s resistance drift in the upwards direction as a result of the impact of $V_{\text{off}}$.
Experiment Max/Min Pulse Pulse Duty
Amplitude Step Duration (s) Cycle
Negative Ascending -0.2 -0.01 1000 0.1
Negative Descending -0.2 -0.01 1000 0.1
Positive Ascending 0.2 0.01 1000 0.1
Positive Descending 0.07 0.01 1000 0.1
Alternating 0.5 N/A 1000 0.001
TABLE III: Parameters for the input signals in the simulated switching experiments. In the case of the positive descending experiment, we observe that too large a signal amplitude drives the memristor immediately to a high resistance state, so we lower the initial amplitude to better observe the behaviour. In the case of the alternating experiment, we use a small duty cycle such that the active pulse duration is 1s.
@@
Negative Ascending Negative Descending Alternating Positive Ascending Positive Descending
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Simulated experiments for a variety of different voltage inputs. We see that the positive/negative pulsing experiments tend to bring about switching in the high/low resistance directions, respectively. In the alternating pulses experiment, the high equilibrium resistance results in the final resistance being higher than the initial resistance, due to resistive drift. Green lines are voltage inputs and red-blue lines are resistance series, with bluer lines indicating a lower initial resistance.
IV Neuromorphic Applications

### v3

@@ IV Metal Oxide Memristor Modelling
In this section, we look at the application of the modelling framework to the modelling of resistive drift phenomenon under zero bias in titanium dioxide memristors. As such, we give an illustrative example of the proposed modelling framework, detailing a rate equation (dictating switching probabilities), the volatility state variables, and a readout equation for this setting.
IV-A Physical Model
@@
Metal oxide Metal-Insulator-Metal (MIM), and more specifically - Titanium Dioxide MIM - has been previously modelled as a single conductive pathway of varying resistance [10], and other similar devices, such as tantalum oxide memristors have been modelled as exhibiting the dominance of a single filament or conductive region [16].
@@
Based on initial investigation of a drift dataset (see Section IV-E below), we model the conductance of the device as linearly proportional to state variable $n$. We hypothesise that the mechanism underlying the switching of the device modulates the width of a conductive region within the active substrate, bridging a connection between the electrodes. Phenomena that could account for this behaviour include a single conductive filament that grows or reduces in width, or multiple conductive filaments that each contribute linearly to the conductance of the device.
@@ IV-B Readout Equation
$R(n)=\frac{1}{g_{\text{diff}}\max{(n-n_{\text{thresh}},0)}+g_{\text{parallel}}}$ (4)
We can see from the equation that the behaviour of the model is thresholded, with the conductance remaining at the value $g_{\text{parallel}}$ until the value of $n$ surpasses the parameter $0\leq n_{\text{thresh}}\leq N$, upon which the conductance increases linearly in steps of $g_{\text{diff}}$. This is motivated by the conductive filament model described above.
@@
$n(R)=n_{\text{thresh}}+\frac{1}{g_{\text{step}}}\left(\frac{1}{R}-g_{\text{parallel}}\right)$ (5)
@@
IV-C Switching Probability
To determine the rates of the Poisson processes, we make use of a Boltzmann occupation function that is dependent on the voltage bias and the temperature, as well as static parameters, rather than the Fermi-Dirac occupation function (logistic function) used to determine the switching probabilities in the GMSM. This is due to it being a good model of energy-barrier modulated processes at thermodynamic equilibrium in classical systems.
@@ IV-C Switching Probability
with the parameters defined in Table I. In this model, application of a voltage bias, $V(t)$, lowers or raises the activation energy barrier in one direction, simultaneously raising or lowering it in the other direction.
$E_{a}=qV_{a}$ Activation energy (barrier height). Increases/decreases the up/down switching rates simultaneously.
$E_{\text{off}}=qV_{\text{off}}$ Energy offset that causes a shift of the equilibrium point from equal switches being in the high/low resistive state. Adjusts the balance of switching rates under zero bias. (Added, rather than subtracted, for $\lambda^{\prime}_{down}$).
@@
$T(t)$ Time-dependent temperature variable in degrees Kelvin. Influenced by the history of applied voltages and the device state history, which directly impact the switching rates, as well as the equilibrium point when $E_{\text{off}}\neq 0$.
$\rho(t)$ Volatility state variable. Has a similar impact to the temperature in affecting the global switching rate (indeed, temperature is a volatility state variable in its own right).
@@
IV-D Volatility
@@
In modelling the effects of Joule heating, we can assume that a proportion of the current flowing through the device (proportional to the square of the current flowing through the device) is converted to heat energy and raises the internal temperature. We can also assume that Newton’s law of cooling applies simultaneously, resulting in power dissipation and restoration of the internal temperature to the equilibrium (bath) temperature. One can define the rate of change of temperature in terms of a thermal time constant that is a function of a thermal resistance ($R_{th}$) and thermal capacitance ($C_{th}$) [17], dictating how easily heat energy in the form of temperature is captured and dissipated from the device, with the thermal time constant denoted $\tau_{th}=C_{th}R_{th}$. The differential equation for the internal temperature of the device is given as:
@@
where $\tau_{\text{volatile}}$ is the volatility time constant, dictating how quickly the equilibrium value is attained (and therefore how slowly a non-zero volatility decays in the absence of an input signal); and $c_{\text{volatile}}$ is the volatility factor, which - when multiplied by the applied bias - yields the equilibrium volatility (dictating the volatility state variable’s leverage on the switching rates). To incorporate the continuous time volatility, we adopt a discrete time sampling approach, with deterministic events at intervals of the given timestep $t_{\text{sample}}=0.1s$. To compute changes to the parameters, we use Euler integration.
IV-E Dataset
@@
IV-E1 Quantisation
Figure 5 shows a histogram of the unquantised resistance values in the dataset, showing the range of interest to be the interval of approximately $10^{3},10^{5}]$, where changes to the value of $n$ significantly affect the output resistance. As such, we choose the values of $N$, $n_{\text{thresh}}$, $g_{\text{step}}$, and $g_{\text{parallel}}$ accordingly. We set $g_{\text{step}}=10^{-7}$, $g_{\text{parallel}}=10^{-10}$, and $N=2n_{\text{thresh}}=20,000$.
@@ IV-E1 Quantisation
$\mathcal{R}_{\text{val}}=$ $\{R(n_{\text{thresh}}),R(n_{\text{thresh}}+1),\cdots,R(N)\}$
where $\begin{cases}|\mathcal{R}_{\text{bound}}|&=N-n_{\text{thresh}}+2,\\ |\mathcal{R}_{\text{val}}|&=N-n_{\text{thresh}}+1.\end{cases}$
IV-E2 Pairs
@@
IV-F Model Fitting
IV-F1 Calculation of the Equilibrium Point (Metastable Point)
@@
$n_{\text{eq}}$ $=\frac{N}{e^{\frac{qV_{\text{off}}}{k_{B}T}}+1}.$ (11)
@@
$V_{\text{off}}$ $=\ln{\left(\frac{N-n_{\text{eq}}}{n_{\text{eq}}}\right)}\cdot\frac{k_{B}T}{q}\equiv\ln{\left(\frac{m_{\text{eq}}}{n_{\text{eq}}}\right)}\cdot V_{T},$ (12)
@@
IV-F2 Linear Conductance Approximation
@@
We use a zero-bias model consisting of an approximately constant change in the conductance over the given time interval, that is independent of the initial resistance. This assumption is based on a linear difference approximation of our readout equation model (Equation 4), which fits a trend in the observed data to a reasonable standard. We use an iterative trust region fitting algorithm [20] to minimise the squared error associated with fitting the function to the predicted resistance change. The model is given as follows, with parameter $a$ - the change in the state variable - fit to the data:
$\Delta R(R_{\text{init}})=\frac{1}{g_{\text{parallel}}-ag_{\text{diff}}}-R_{\text{init}}.$ (13)
For the chosen timestep of $10,000$ seconds, we find a value of $a=3.62$ yields the best fit, for the given parameters $g_{\text{diff}}$, $g_{\text{parallel}}$, $N$, and $n_{\text{thresh}}$. This associated line of best fit for this linear conductance change model is shown in Figure 6. The moving averages are plotted to enable a clearer view of the general trend. Our linear conductance model is also accurate across different values of $\Delta T$, as shown in Figure 7. We choose $\Delta T=10000$ since larger values of $\Delta T$ result in datasets with higher signal to noise ratios, since stochastic fluctuations in the memristive state will become smaller compared to the deterministic drift component, allowing for better estimation of the associated model parameter, $a$.
Figure 6: The mean change in resistance for the dataset of pairs with $\Delta T=10000$. We show a moving average of the data for different window sizes (5 and 50), to see the non-linear trend.
Figure 7: Fitting the parameter $a$ of the linear conductance model for different timesteps, $\Delta T$, showing the estimate is proportional to the timestep, supporting the linear assumption.
IV-F3 Parameter Fitting and Constant Rate Approximation
@@
We would like to obtain parameters that allow the mean behaviour of the model to correspond to the mean predictions of the linear conductance model. Due to the difficulty in obtaining an analytic expression from a finite dataset for the delay-conditional distribution $P(R(t+\Delta T)|R(t),\Delta T)$ without sufficient knowledge of the nature of the underlying distribution, we use a constant rate approximation to obtain approximate mean values for this distribution over time. We assume that the up and down switching rates of the model remain relatively constant. In order to fit the model, we calculate the rate at a value of $n=N$ and make the assumption that the rate does not vary greatly over the region of interest.
@@
As such, we require $N>>a$. We must also ensure that $N+1-n_{\text{thresh}}$, defining the number of unique resistance states according to Equation 4, is appropriately selected to give a sufficiently fine quantisation over the switching region for which we have data.
@@
For the given value of the parameter $a$, we can determine an appropriate value of $V_{a}$. We first select a value of $V_{\text{off}}$ to ensure an equilibrium point (see Section IV-F1) that is far above the active switching region (in terms of resistance), to ensure that the assumption that the rate is constant over the considered range holds, since the equilibrium point at a particular bias sets the zero-crossing point for the rate equation, hence our constant rate assumption will break down as the value of $n$ approaches $n_{\text{eq}}$. In our case, this corresponds to setting an equilibrium point with a value of $n_{\text{eq}}<<n_{\text{thresh}}$. We can see in Figure 8 how the choice of $V_{\text{off}}$ impacts $n_{\text{eq}}$; and hence, the constant rate assumption, and the fit to the linear conductance model. Assuming a constant rate and a given value of $V_{\text{off}}$, we can derive a value of $V_{a}$ according to the following equation:
$V_{a}=\frac{k_{B}T}{q}\ln{\left(2\cdot\Delta T\cdot N\cdot\sinh{\left(V_{\text{off}}\frac{q}{2k_{B}T}\right)/a}\right)}.$ (14)
@@
In Figure 9, we see averaged zero-bias simulations from the event-based model, based on the fit parameters, showing the mean change in the resistance fitting to the linear conductance model, as well as the associated variance of the evolutions.
Figure 9: The mean change in the resistance following fitting to the linear conductance model as compared to the data and the linear conductance model, with confidence bands indicating the standard deviation of the data and the simulated series, for simulations with length $\Delta T=10000$.
IV-G Simulation
In addition to demonstrating the fit of the zero-bias mean values to those determined according to the linear conductance model, we demonstrate the capabilities of the model to simulate memristive behaviour through a series of simulated experiments for a titanium dioxide thin film memristive device. Due to the event-driven nature of the state transition model we conduct our simulations in this section and in Section VI in Python, rather than a circuit simulation language such as SPICE.
IV-G1 Experimental Parameters
The parameters used for the experiments, chosen for $\Delta T=10000$, are summarised in Table II. The choices for these parameters are given in Sections IV-E1 and IV-F3.
@@ IV-G1 Experimental Parameters
IV-G2 Switching Experiments
@@
Negative Ascending Alternating Positive Ascending
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Simulated experiments different voltage inputs. We see that the positive/negative pulses switch in the high/low resistance directions, respectively. The high equilibrium resistance results in the final resistance being higher than the initial resistance.
V Model Accuracy and Efficiency
Here, we evaluate aspects of the accuracy and the computational efficiency of the model, in the context of its applicability to neuromorphic computing.
V-A Impact of Parameter NN
One strength of a discrete modelling strategy is that simulations can be conducted at different resolutions, depending on precision requirements, and the scale of interest. For example, if a memristive device is to be used as a binary storage device, where only the states $1$ and 0 are distinct, then a value of $N=2$ suffices to model the readout value of interest. In contrast, for a simulation where multiple stable states are to be for information storage, a larger value of $N$ may be required. As a consequence of the fact that the variance and mean of a Poisson process are equal to the rate of the process, the magnitude of the parameter $N$ will scale the expected value and the variance of the final state in a linear fashion. We can retain the variance and mean by adjusting the parameter $V_{\text{a}}$ to offset the impact of $N$ on the simulation statistics. In the case of aligning our model to the phenomenon of drift, as previously shown in Figure 9, we can use Equation 14 to set $V_{\text{a}}$ to keep the rate parameter the same, for a given $V_{\text{off}}$ and updated value of $N$. In Figure 12(a), we compare resistive drift simulations for different values of $N$, with $V_{a}$ adjusted to achieve the same mean and variance (rate parameter). Runtimes vs. errors for the simulations are shown in Figure 12(c).
V-B Discrete-Time Sampling Comparison
We note that the approximation in Equation 1 is valid only for small values of the discrete timestep $h$. If $h$ grows large, then discretisation errors emerge. The larger $h$ becomes, the more significant a change will have occurred in the parameters used to calculate the transition probabilities, leading to inaccuracies. An advantage of our event-based adaptation of the GMSM is to eliminate errors associated with discrete time-stepped simulation. However, it should be noted that using a value of N that is too coarse will still introduce a quantisation error.
A discrete time Euler approximation of the event-based switching model can be achieved by using a normal approximation to the number of events generated by a Poisson process over a given timestep $h$ as $\mathcal{N}(\lambda h,\lambda h)$. Figures 12(b) and 12(c) show 100 simulations per timestep of the discrete time switching approximation and a corresponding error-runtime curve, demonstrating the trade-off between decreasing the timestep size for accuracy and increasing the computational burden. We see that a discrete time approximation corresponds to the event-based switching to a reasonable degree, and this approach may offer benefits in settings where the number of switching events per unit time is high. In our case, we limit our modelling to the resistive drift setting, for which the number of switching events is low, leading to the event-based model outperforming the discrete time model in both accuracy and runtime. In the case of incorporation of event-based approximations of continuous parameter updates, such as those for volatility state variables, as outlined in Section III-D3, we can upper bound the error of the event-based model with a discrete time modelling error for a timestep $h$. This frames the event-based modelling approach as an extension of discrete time modelling, with additional parameter updates occurring dynamically where they are prompted by model changes.
(a) Simulations for different resolutions (NN values), with $V_{\text{a}}$ adjusted to give equivalence.
(b) Discrete time approximations of the event-based simulation, for increasing $h$.
(c) Runtimes vs. average errors for the discrete and event-based modelling approaches.
Figure 12: Comparison of a) event-based for different NN, and b) discrete timestep based simulation for different $h$, with runtime vs. error shown in c). In c), each event-based simulation has the same runtime due to the Poisson rates being the same. We compute the error as the sum of the relative square errors in the mean and variance averaged over initial resistances.
Finally, we note that circuit simulation software, such as SPICE, is also capable of implementing adaptive timesteps for input signals, which may offer a good tradeoff between accuracy and speed. However, for simulations involving stochastic state evolution, such approaches are less suitable, as they rely on an approximation of the instantaneous rate of change of parameters to compute the required timestep size, which is not straightforward to estimate for stochastic dynamics.
VI Neuromorphic Applications